# Retrieval-Augmented Generation for AI-Generated Content: A Survey

### v5

@@ Abstract
Advancements in model algorithms, the growth of foundational models, and access to high-quality datasets have propelled the evolution of Artificial Intelligence Generated Content (AIGC). Despite its notable successes, AIGC still faces hurdles such as updating knowledge, handling long-tail data, mitigating data leakage, and managing high training and inference costs. Retrieval-Augmented Generation (RAG) has recently emerged as a paradigm to address such challenges. In particular, RAG introduces the information retrieval process, which enhances the generation process by retrieving relevant objects from available data stores, leading to higher accuracy and better robustness. In this paper, we comprehensively review existing efforts that integrate RAG technique into AIGC scenarios. We first classify RAG foundations according to how the retriever augments the generator, distilling the fundamental abstractions of the augmentation methodologies for various retrievers and generators. This unified perspective encompasses all RAG scenarios, illuminating advancements and pivotal technologies that help with potential future progress. We also summarize additional enhancements methods for RAG, facilitating effective engineering and implementation of RAG systems. Then from another view, we survey on practical applications of RAG across different modalities and tasks, offering valuable references for researchers and practitioners. Furthermore, we introduce the benchmarks for RAG, discuss the limitations of current RAG systems, and suggest potential directions for future research. Github: https://github.com/PKU-DAIR/RAG-Survey.
@@ I-A Background
Despite the remarkable progress made by advanced generative models, AIGC continues to face a number of well-known challenges, including the struggle to maintain up-to-date knowledge, the inability to incorporate long-tail knowledge [27], and the risk of leaking private training data [28]. Retrieval-Augmented Generation (RAG) is proposed to alleviate, if not completely address, the aforementioned challenges through its adaptable data repository [29]. The knowledge stored for retrieval can be conceptualized as non-parametric memory, which is easily modifiable, capable of accommodating broad long-tail knowledge, and also able to encode confidential data. In addition, retrieval can also be employed to reduce the generation costs. For example, RAG can reduce the size of large generative models [30], provide support for long contexts [31], and eliminate certain generation steps [32].
A typical RAG process is depicted in Fig. 1. Given an input query, the retriever locates and looks up relevant data sources, then the retrieved results interact with the generator to enhance the overall generation process. There are several foundational paradigms (foundations in short) according to how the retrieved results augment the generation: they can serve as augmented input to the generator [33, 34]; they can join at the middle stage of generation as latent representations [35, 36]; they can contribute to the final generation results in the form of logits [37, 38]; they can even influence or omit certain generation steps [32, 39]. Moreover, beyond the foundational RAG process, researchers have proposed numerous enhancements to elevate the overall quality. These methods encompass specific optimizations for individual components as well as holistic enhancements aimed at the entire pipeline.
@@ I-B Contribution
This survey offers a comprehensive overview of RAG, covering foundations, enhancements, applications, benchmarks, limitations, and potential future directions. While retrievers and generators exhibit variations across modalities and tasks, we distill the fundamental abstractions of RAG foundations, considering applications as adaptations stemming from these abstractions. We aim to offer references and guidelines to researchers and practitioners, providing valuable insights for advancing RAG methodologies and related applications. In summary, we list our contributions as follows:
@@ I-C Related Work
In addition, another survey proposed by Zhao et al. [60] introduces RAG applications across multiple modalities, but ignoring the discussion on RAG foundations. While existing research has explored various aspects of RAG, there remains a need for a comprehensive overview that covers RAG foundations, enhancements, and its applicability across different domains. In this paper, we aim to address the gap by presenting a systematic survey of RAG.
@@ II-A1 Transformer Model
II-A2 LSTM
As shown in Fig. 2, Long Short-Term Memory (LSTM) [63] is a special form of Recurrent Neural Network (RNN) model. It addresses the exploding/vanishing gradient problems when processing long-term dependency information by introducing cell state and gate mechanisms. The model consists of three gates (Input Gate, Forget Gate, and Output Gate), which serve as filters of information, and a core module Cell State, which can memorize and maintain information. It uses the same vocabulary classification method as transformer models to autoregressively generate outputs.
II-A3 Diffusion Model
@@ II-B2 Dense Retriever
During training, dense retrieval usually follows a contrastive learning paradigm, making positive samples more similar and negative samples less similar. Several hard negative techniques [70] have been proposed to further improve the model quality. During inference, approximate nearest neighbor (ANN) methods are applied for efficient searching. Various indices are developed to serve ANN search, such as tree [71, 72], locality sensitive hashing [73], neighbor graph index (e.g., HNSW [74], DiskANN [75]), and the combination of graph index and inverted index (e.g., SPANN [22]).
II-B3 Others
In addition to sparse retrieval and dense retrieval, there are alternative methods for retrieving relevant objects [76, 77]. Instead of calculating representations, some research works directly use the edit distance between natural language texts [78] or abstract syntax trees (AST) of code snippets [79, 80]. For knowledge graph, entities are linked with relations, which can be regarded as a pre-built index for retrieval searching. Therefore, RAG methods which involve knowledge graph can use $k$-hop neighbor search as retrieval process [81, 82]. Named entity recognition (NER) [83] is another way of retrieval, where the input is the query and the entites are the keys.
III Methods
In this section, we introduce RAG foundations and outline enhancement methods that further improve the effectiveness.
@@ III-A1 Query-based RAG
Stemming from the idea of prompt augmentation, Query-based RAG seamlessly integrates the user’s query with insights from retrieved information, feeding it directly into the initial stage of the language model’s input. This method is prevalent in RAG applications. Post-retrieval, the obtained content is merged with the user’s original query to form a composite input sequence, which is then processed by the generator to create a response. Query-based RAG is widely employed across various modalities.
@@
In the task of Image Generation, RetrieveGAN [45] enhances the relevance and accuracy of generated images by integrating retrieved information, including selected image patches and their corresponding bounding boxes, into the input stage of the generator. IC-GAN [94] modulates the specific conditions and details of the generated images by concatenating noise vectors with instance features.
@@ III-A2 Latent Representation-based RAG
In the Text field, FiD [35] and RETRO [36] are two classic structures of Latent Representation-based RAG, with many subsequent works being modifications based on them. FiD [35] processes each retrieved paragraph and its title alongside the query through distinct encoders, then amalgamates the resulting latent representations for decoding by a single decoder to produce the final output. RETRO [36] retrieves relevant information for each segmented sub-query, then applies a novel module termed Chunked Cross-Attention (CCA) to integrate the retrieved contents with each sub-query tokens. In addition, there are other noteworthy novel structures within the scope of latent representation-based RAG. Several studies [31, 95] have integrated kNN search within transformer blocks, allowing for input chunking and, in theory, addressing the long-criticized context length constraints of Transformer models. Kuratov et al. [96] integrated Transformer with RNN, utilizing the model’s intermediate output as the content for retrieval.
In the realms of Code, and Science, FiD has gained widespread adoption, with applications spanning code-related fields [97, 98, 99, 100, 101], and AI-for-Science [55], among others.
@@
For video captioning, R-ConvED [48] uses a convolutional encoder-decoder network to process retrieved video-sentence pairs with an attention mechanism, generating hidden states to produce captions. CARE [116] introduces a concept detector to produce concept probabilities, and incorporates concept representations into a hybrid attention mechanism. EgoInstructor [49] uses gated-cross attention to merge text and video features, improving the relevance and coherence of captions for egocentric videos.
Latent Representation-based RAG is versatile across modalities and tasks, merging retriever and generator hidden states, though it needs extra training for latent space alignment. This approach allows for the creation of complex, innovative algorithms that efficiently integrate retrieved information.
@@ III-A3 Logit-based RAG
In the text domain, kNN-LM [37] and its variant [38] combine language model-generated probabilities with those derived from retrieval distances of similar prefixes at each decoding step. TRIME [117] and NPM [118] are radical evolutions of traditional kNN-LM approaches, using closely aligned tokens from a local database as output, particularly boosting performance in long-tail distribution scenarios.
Beyond Text, other modalities, such as code and Image, also leverage logit-based RAG. In the domain of code, several studies [79, 119] have also adopted the concept KNN to enhance final output control, thereby achieving superior performance. Furthermore, EDITSUM [97] improves the quality of code summarization by integrating prototype summaries at the logit level. For Image captioning, MA [120] directly applies the KNN-LM framework to address the image caption problem, achieving favorable results.
In summary, Logit-based RAG utilizes historical data to deduce current states and merges information at the logit level, ideal for sequence generation. It focuses on generator training and allows for novel methods that capitalize on probability distributions for future tasks.
@@ III-B1 Input Enhancement
Query Transformation
Query transformation can enhance the result of retrieval by modifying the input query.
Query2doc [124] and HyDE [125] use the original query to generate a pseudo document, which is later used as the query for retrieval. The pseudo document contains richer relevant information, which helps to retrieve more accurate results. TOC [126] leverages retrieved contents to decompose the ambiguous query into multiple clear sub-queries, which are sent to the generator and aggregated to produce the final result.
Data Augmentation
Data augmentation improves data before retrieval, including techniques such as removing irrelevant information, eliminating ambiguity, updating outdated documents, synthesize new data, etc.
Make-An-Audio [44] uses captioning and audio-text retrieval to generate captions for language-free audio to mitigate data sparsity, and adds random concept audio to improve the original audio. LESS [127] strategically selects an optimal dataset for downstream tasks by analyzing gradient information. It aims to maximize the dataset’s impact on fine-tuning the model’s performance in response to instructional prompts. ReACC [90] employs data augmentation (including renaming and dead code insertion) to pre-train the code retrieval model.
@@ III-B2 Retriever Enhancement
Recursive Retrieval
Recursive retrieval is to perform multiple searches to retrieve richer and higher-quality contents.
@@
Chunk optimization refers to adjusting chunk size for improved retrieval results.
@@
LlamaIndex [131] incorporates a series of chunk optimization methods, one of which operates on a ’small to big’ principle. The core concept here is to pinpoint finer-grained content but return richer information. For instance, Sentence-window retrieval fetches small text chunks and returns a window of relevant sentences surrounding the retrieved segment. In Auto-merge retrieval, documents are arranged in a tree structure. The process retrieves the parent node, which encapsulates the content of its child nodes, by fetching the child node first. To address the lack of contextual information, RAPTOR [132] employs recursive embedding, clustering, and summarization of text chunks until further clustering becomes infeasible, thereby constructing a multi-level tree structure.
Retriever Finetuning
@@
The retriever, central to the RAG system, relies on a proficient embedding model [133, 134, 135, 136] to cluster related content and feed the generator, enhancing system performance.
@@
Additionally, embedding models with strong expressive power can be fine-tuned with domain-specific or task-related data to boost performance in targeted areas. REPLUG [85] treats LM as a black box and update the retriever model based on the final results. APICoder [87] finetunes the retriever with python files and api names, signature, description. EDITSUM [97] finetunes the retriever to decrease the jaccard distance between summaries after retrieval. SYNCHROMESH [80] adds tree distance os ASTs in the loss and uses Target Similarity Tuning to finetune the Retriever. R-ConvED [48] finetunes the Retriever with the same data as generator. Kulkarni et al. [137] applied infoNCE loss to finetune the Retriever.
Hybrid Retrieval
@@
Hybrid retrieve denotes the concurrent employment of a diverse array of retrieval methodologies or the extraction of information from multiple distinct sources.
@@
RAP-Gen [138] and ReACC [90] use both dense retriever and sparse retriever to improve the quality of retrieval. Rencos [79] uses sparse retriever to retrieve similar code snippets on syntactic-level and usse dense retriever to retrieve similar code snippets on semantic-level. BASHEXPLAINER [98] first uses dense retriever to capture semantic information and then uses sparse retriever to acquire lexical information. RetDream [50] first retrieves with text and then retrieves with the image embedding. CRAG [139] features a retrieval evaluator that gauges document relevance to queries, prompting three retrieval responses based on confidence: direct use of results for Knowledge Refinement if accurate, Web Search if incorrect, and a hybrid approach for ambiguous cases. Huang et al. [140] improved question-answering by introducing DKS (Dense Knowledge Similarity) and RAC (Retriever as Answer Classifier) in the retrieval phase, evaluating answer relevance and knowledge applicability. UniMS-RAG [141] introduces a novel kind of token, termed as the “acting token”, which determines the source from which to retrieve information.
Re-ranking
@@
The Rerank technique refers to reordering the retrieved content in order to achieve greater diversity and better results. Re2G [142] applies a re-ranker [143] model after the traditional retriever to reduce the impact of information loss caused by compressing text into vectors. AceCoder [144] reranks the retrieved programs with a selector to reduce redundant programs and obtain diverse retrieved programs. XRICL [145] uses a distillation-based exemplar reranker after retrieval. Rangan [146] employs the Quantized Influence Measure, assessing statistical biases between a query and a reference to evaluate the similarity of data subsets and rerank retrieval results. UDAPDR [147] uses LLMs to cost-effectively generate synthetic queries that train domain-specific rerankers, which then apply multi-teacher knowledge distillation to develop a cohesive retriever. LLM-R [148] refines its retriever iteratively by employing a static LLM for document ranking and reward model training, complemented by knowledge distillation. Each training cycle incrementally improves the retriever, enabling progressive optimization.
Retrieval Transformation
@@
Retrieval Transformation involves rephrasing retrieved content to better activate the generator’s potential, resulting in improved output.
@@
Others
In addition to the above optimization methods, there are also some other optimization methods for the retrieve process. For example, Meta-data filtering  [152] is a method to help processing retrieved documents which uses metadata (such as time, purpose, etc.) to filter the retrieved documents for better results. GENREAD [153] and GRG [154] introduce a novel approach where the retrieval process is supplanted or improved by prompting a LLM to generate documents in response to a given question.

### v6

@@ Abstract
Advancements in model algorithms, the growth of foundational models, and access to high-quality datasets have propelled the evolution of Artificial Intelligence Generated Content (AIGC). Despite its notable successes, AIGC still faces hurdles such as updating knowledge, handling long-tail data, mitigating data leakage, and managing high training and inference costs. Retrieval-Augmented Generation (RAG) has recently emerged as a paradigm to address such challenges. In particular, RAG introduces the information retrieval process, which enhances the generation process by retrieving relevant objects from available data stores, leading to higher accuracy and better robustness. In this paper, we comprehensively review existing efforts that integrate RAG techniques into AIGC scenarios. We first classify RAG foundations according to how the retriever augments the generator, distilling the fundamental abstractions of the augmentation methodologies for various retrievers and generators. This unified perspective encompasses all RAG scenarios, illuminating advancements and pivotal technologies that help with potential future progress. We also summarize additional enhancements methods for RAG, facilitating effective engineering and implementation of RAG systems. Then from another view, we survey on practical applications of RAG across different modalities and tasks, offering valuable references for researchers and practitioners. Furthermore, we introduce the benchmarks for RAG, discuss the limitations of current RAG systems, and suggest potential directions for future research. Github: https://github.com/PKU-DAIR/RAG-Survey.
@@ I-A Background
Despite significant advancements in generative models, AIGC still grapples with challenges like outdated knowledge, lack of long-tail knowledge [27], and risks of leaking private training data [28]. Retrieval-Augmented Generation (RAG) aims to mitigate these issues with its flexible data repository [29]. The retrievable knowledge acts as non-parametric memory, which is easily updatable, accommodates extensive long-tail knowledge, and can encode confidential data. Moreover, retrieval can lower generation costs. RAG can reduce the size of large models [30], support long contexts [31], and eliminate certain generation steps [32].
A typical RAG process is depicted in Fig. 1. Given an input query, the retriever identifies relevant data sources, and the retrieved information interacts with the generator to improve the generation process. There are several foundational paradigms (foundations in short) according to how the retrieved results augment the generation: they can serve as augmented input to the generator [33, 34]; they can join at the middle stage of generation as latent representations [35, 36]; they can contribute to the final generation results in the form of logits [37, 38]; they can even influence or omit certain generation steps [32, 39]. Additionally, researchers have proposed various enhancements to improve the foundational RAG process. These methods encompass specific optimizations for individual components as well as holistic enhancements aimed at the entire pipeline.
@@ I-B Contribution
This survey offers a comprehensive overview of RAG, covering foundations, enhancements, applications, benchmarks, limitations, and potential future directions. Despite variations in retrievers and generators across modalities and tasks, we distill the core principles of RAG foundations, viewing applications as adaptations of these principles. We aim to offer references and guidelines to researchers and practitioners, providing valuable insights for advancing RAG methodologies and related applications. In summary, we list our contributions as follows:
@@ I-C Related Work
In addition, another survey proposed by Zhao et al. [60] introduces RAG applications across multiple modalities, but ignoring the discussion on RAG foundations. Another work [61] covers only part works of other modalities. While existing research has explored various aspects of RAG, there remains a need for a comprehensive overview that covers RAG foundations, enhancements, and its applicability across different domains. In this paper, we aim to address the gap by presenting a systematic survey of RAG.
@@ II-B1 Transformer Model
II-B2 LSTM
As shown in Fig. 2, Long Short-Term Memory (LSTM) [64] is a special form of Recurrent Neural Network (RNN) model. It tackles the issues of exploding/vanishing gradients in long-term dependency processing by incorporating cell states and gating mechanisms. The model comprises three gates (Input, Forget, and Output) that filter information, and a central Cell State module that retains and manages information. It uses the same vocabulary classification method as transformer models to autoregressively generate outputs.
II-B3 Diffusion Model
@@ II-C2 Dense Retriever
During training, dense retrieval uses contrastive learning to increase the similarity of positive samples and decrease that of negative ones. Several hard negative techniques [71] have been proposed to further enhance model quality. For efficient searching during inference, ANN methods are employed. Various indices are developed to serve ANN search, including tree [72, 73], locality sensitive hashing [74], neighbor graph indices (e.g., HNSW [75], DiskANN [76]), and combined graph and inverted indices (e.g., SPANN [22]).
II-C3 Others
In addition to sparse retrieval and dense retrieval, there are alternative methods for retrieving relevant objects [77, 78]. Instead of calculating representations, some research works directly use the edit distance between natural language texts [79] or abstract syntax trees (AST) of code snippets [80, 81]. In knowledge graphs, entities are connected by relations, serving as a pre-built index for retrieval. Thus, RAG methods utilizing knowledge graphs can employ $k$-hop neighbor searches for retrieval [82, 83]. Another retrieval method is Named Entity Recognition (NER) [84], where the query is the input and the entities act as keys.
III Methodologies
In this section, we first introduce foundational paradigms of RAG, and then outline enhancement methods that further improve the effectiveness.
@@ III-A1 Query-based RAG
Stemming from the idea of prompt augmentation, query-based RAG seamlessly integrates the user’s query with insights from retrieved information, feeding it directly into the initial stage of the generator’s input. This method is prevalent in RAG applications. Post-retrieval, the obtained content is merged with the user’s original query to form a composite input, which is then processed by the generator to create a response. Query-based RAG is widely employed across various modalities.
@@
In the image generation task, RetrieveGAN [45] boosts the relevance and precision of generated images by incorporating retrieved data, such as selected image patches and their bounding boxes, into the generator’s input stage. IC-GAN [95] modulates the specific conditions and details of the generated images by concatenating noise vectors with instance features.
@@ III-A2 Latent Representation-based RAG
In the text field, FiD [35] and RETRO [36] are two classic structures of latent representation-based RAG, with many subsequent works conducting modifications based on them. FiD [35] processes each retrieved paragraph and its title alongside the query through distinct encoders, then amalgamates the resulting latent representations for decoding by a single decoder to produce the final output. RETRO [36] retrieves relevant information for each segmented sub-query, then applies a novel module termed Chunked Cross-Attention (CCA) to integrate the retrieved contents with each sub-query tokens. In addition, there are other noteworthy novel structures within the scope of latent representation-based RAG. Several studies [31, 96] have integrated k Nearest Neighbor (kNN) search within transformer blocks, allowing for input chunking and, in theory, addressing the long-criticized context length constraints of Transformer models. Kuratov et al. [97] integrated Transformer with RNN, utilizing the model’s intermediate output as the content for retrieval.
In the realms of code and science, FiD has gained widespread adoption, with applications spanning various code-related fields [98, 99, 100, 101, 102], and AI-for-Science [55].
@@
For video captioning, R-ConvED [48] uses a convolutional encoder-decoder network to process retrieved video-sentence pairs with an attention mechanism, generating hidden states to produce captions. CARE [117] introduces a concept detector to produce concept probabilities, and incorporates concept representations into a hybrid attention mechanism.
EgoInstructor [49] uses gated-cross attention to merge text and video features, improving the relevance and coherence of captions for egocentric videos. Latent representation-based RAG, adaptable across modalities and tasks, blends retriever and generator hidden states but requires additional training for aligning latent spaces. It enables the development of sophisticated algorithms that seamlessly incorporate retrieved information.
@@ III-A3 Logit-based RAG
In the text domain, kNN-LM [37] and its variant [38] blend language model probabilities with those from retrieval distances of similar prefixes at each decoding step. TRIME [118] and NPM [119] are radical evolutions of traditional kNN-LM approaches, using closely aligned tokens from a local database as output, particularly boosting performance in long-tail distribution scenarios.
Beyond text, other modalities, such as code and image, also leverage logit-based RAG.
In the domain of code, several studies [80, 120] have also adopted the concept kNN to enhance final output control, thereby achieving superior performance. Furthermore, EDITSUM [98] improves the quality of code summarization by integrating prototype summaries at the logit level. For image captioning, MA [121] directly applies the kNN-LM framework to address the image caption problem, achieving favorable results. In summary, logit-based RAG utilizes historical data to deduce current states and merges information at the logit level, ideal for sequence generation. It focuses on generator training and allows for novel methods that capitalize on probability distributions for future tasks.
@@ III-B1 Input Enhancement
Query Transformation: Query transformation can enhance the result of retrieval by modifying the input query.
Query2doc [125] and HyDE [126] use the original query to generate a pseudo document, which is later used as the query for retrieval. The pseudo document contains richer relevant information, which helps to retrieve more accurate results. TOC [127] leverages retrieved contents to decompose the ambiguous query into multiple clear sub-queries, which are sent to the generator and aggregated to produce the final result. For complex or ambiguous queries, RQ-RAG [128] breaks them down into clear subqueries for fine-grained retrieval and synthesizes the responses to deliver a cohesive answer to the original query. Tayal et al. [129] refined the initial query using dynamic few-shot examples and context retrieval, enhancing the generator’s grasp of user intent.
Data Augmentation: Data augmentation improves data before retrieval, including techniques such as removing irrelevant information, eliminating ambiguity, updating outdated documents, synthesize new data, etc.
Make-An-Audio [44] uses captioning and audio-text retrieval to generate captions for language-free audio to mitigate data sparsity, and adds random concept audio to improve the original audio. LESS [130] optimizes dataset selection for downstream tasks by analyzing gradient information, aiming to enhance model performance in response to instructional prompts. ReACC [91] employs data augmentation (including renaming and dead code insertion) to pre-train the code retrieval model. Telco-RAG [131] enhances the retrieve accuracy by appling a “Vocabulary for 3GPP Specifications”, and match them to user queries with a router module.
@@ III-B2 Retriever Enhancement
Recursive Retrieval: Recursive retrieval is to perform multiple searches to retrieve richer and higher-quality contents.
@@
Chunk Optimization: Chunk optimization refers to adjusting chunk size for improved retrieval results.
@@
LlamaIndex [135] incorporates a series of chunk optimization methods, one of which operates on a ‘small to big’ principle. The core concept here is to pinpoint finer-grained content but return richer information. For instance, Sentence-window retrieval fetches small text chunks and returns a window of relevant sentences surrounding the retrieved segment. In auto-merge retrieval, documents are arranged in a tree structure. The process retrieves the parent node, which encapsulates the content of its child nodes, by fetching the child node first. To address the lack of contextual information, RAPTOR [136] employs recursive embedding, clustering, and summarization of text chunks until further clustering becomes infeasible, thereby constructing a multi-level tree structure. Prompt-RAG [137] enhances retrieval accuracy by pre-generating a table of contents, enabling the model to autonomously select relevant chapters based on the query. Raina et al. [138] break text chunks into finer atomic statements to achieve higher recall and improved results.
@@
Retriever Finetuning: The retriever, central to the RAG system, relies on a proficient embedding model [139, 140, 141, 142] to represent related content and feed the generator, enhancing system performance.
@@
Additionally, embedding models with strong expressive power can be fine-tuned with domain-specific or task-related data to boost performance in targeted areas. REPLUG [86] treats LM as a black box and update the retriever model based on the final results. APICoder [88] finetunes the retriever with python files and api names, signature, description. EDITSUM [98] finetunes the retriever to decrease the jaccard distance between summaries after retrieval. SYNCHROMESH [81] adds tree distance os ASTs in the loss and uses Target Similarity Tuning (TST) to finetune the retriever. R-ConvED [48] finetunes the retriever with the same data as generator. Kulkarni et al. [143] applied infoNCE loss to finetune the retriever.
@@
Hybrid Retrieval: Hybrid retrieve denotes the concurrent employment of a diverse array of retrieval methodologies or the extraction of information from multiple distinct sources.
@@
RAP-Gen [144], BlendedRAG [145]and ReACC [91] use both dense retriever and sparse retriever to improve the quality of retrieval. Rencos [80] uses sparse retriever to retrieve similar code snippets on syntactic-level and uses dense retriever to retrieve similar code snippets on semantic-level. BASHEXPLAINER [99] first uses dense retriever to capture semantic information and then uses sparse retriever to acquire lexical information. RetDream [50] first retrieves with text and then retrieves with the image embedding. CRAG [146] features a retrieval evaluator that gauges document relevance to queries, prompting three retrieval responses based on confidence: direct use of results for Knowledge Refinement if accurate, Web Search if incorrect, and a hybrid approach for ambiguous cases. Huang et al. [147] improved question-answering by introducing DKS (Dense Knowledge Similarity) and RAC (Retriever as Answer Classifier) in the retrieval phase, evaluating answer relevance and knowledge applicability. UniMS-RAG [148] introduces a novel kind of token, termed as the “acting token”, which determines the source from which to retrieve information. Koley et al. [149] enhance image retrieval by integrating sketch and text for fine-grained retrieval, yielding improved results.
Re-ranking: The Rerank technique refers to reordering the retrieved content in order to achieve greater diversity and better results.
@@
Re2G [150] applies a re-ranker [151] model after the traditional retriever to reduce the impact of information loss caused by compressing text into vectors. AceCoder [152] reranks the retrieved programs with a selector to reduce redundant programs and obtain diverse retrieved programs. XRICL [153] uses a distillation-based exemplar reranker after retrieval. Rangan [154] employs the Quantized Influence Measure, assessing statistical biases between a query and a reference to evaluate the similarity of data subsets and rerank retrieval results. UDAPDR [155] uses LLMs to cost-effectively generate synthetic queries that train domain-specific rerankers, which then apply multi-teacher knowledge distillation to develop a cohesive retriever. LLM-R [156] refines its retriever iteratively by employing a static LLM for document ranking and reward model training, complemented by knowledge distillation. Each training cycle incrementally improves the retriever, enabling progressive optimization. Finardi et al. [157] integrated reciprocal rank into the retrieval process for enhanced text chunk relevance, and utilized monoT5 as a reranker to optimize the result quality. Li et al. [158] integrate a reranking module into their end-to-end RAG system, enhancing the retrieval quality and factual accuracy of LLMs.
@@
Retrieval Transformation: Retrieval Transformation involves rephrasing retrieved content to better activate the generator’s potential, resulting in improved output.
@@
Others: In addition to the above optimization methods, there are also some other optimization methods for the retrieve process.
For example, meta-data filtering  [162] is a method to help processing retrieved documents which uses metadata (such as time, purpose, etc.) to filter the retrieved documents for better results. GENREAD [163] and GRG [164] introduce a novel approach where the retrieval process is supplanted or improved by prompting a LLM to generate documents in response to a given question. Multi-Head-RAG [165] employs multiple embedding models to project the same text chunk into various vector spaces and utilizes a multi-head attention layer to capture different informational aspects, thereby increasing the accuracy of the retrieval process.